OpenStudy (anonymous):

Find x in this proportion. x =3 Correct ?

OpenStudy (katrinakaif):

Where is your proportion?

OpenStudy (anonymous):



OpenStudy (anonymous):

yeah your right

OpenStudy (anonymous):

good job

OpenStudy (katrinakaif):

Cross Multiply: 5x = 3(x+2) 5x = 3x + 6 Solve 5x - 3x = 6 2x = 6 x= 3 Good job